OpenStudy (anonymous):

An unbiased coin is tossed four times. Find the probability of the given event. (Round your answer to four decimal places.) The coin lands heads more than once.

OpenStudy (blues):

The probability of the coin landing on head with each toss is 1/2. The probability of two or more independent events all happening is the probability of the individual events, multiplied. So where X = the number of heads in four coin tosses: P(X = 2) = 0.5 ^ 2 = 0.25 P(X = 3) = 0.5 ^3 = 0.125 P(X = 4) = 0.5 ^4 = 0.0625 So the probability of all those things happening is P(X = 2) + P(X = 3) + P(X = 4). 0.25 + 0.125 + 0.0625 = 0.4375. I'm a biologist so check my computation.